How many ounces of ceylon tea worth 1.50 per ounce and how many ounces of formosa tea worth 2.00 per ounce must be mixed to obtain a mixture of 8 ounces that is worth 1.85 per ounce..

Please include a chart

Answer by stanbon(75887)

How many ounces of ceylon tea worth 1.50 per ounce and how many ounces of formosa tea worth 2.00 per ounce must be mixed to obtain a mixture of 8 ounces that is worth 1.85 per ounce..
--------
Equation:
1.5c + 2f = 1.85*8
c + f = 8
-------------------
Modify:
1.5c + 2f = 1.85*8
2c + 2f = 16
---------------------
Subtract and solve for "c":
0.5c = 1.2
c = 2.4 ounces (amt. of ceylon tea needed)
Solve for "f":: c + f = 8, so f = 5.2 ounces (amt. of formosa tea needed)
--------------
Cheers,
Stan H.
---------------